A rectangular garden is 15 by 12 meters, you need to have a rectangular area inside such that a uniform strip of side walk is made, the area inside has to be 108 square meters. Find the dimensions of the interior area.
Answer by Paul(988)

108=(15+2x)(12+2x)-(15*20)
108=180+4x^2+54x-180
4x^2+54x-108=0
x^2+13.5x-27=0
solve that you get
x=1.77
Hence, the uniform side is 1.77cm
Paul.